OpenStudy (linstead_cpd_freak):

Solve. 2x + 6 ≥ –4 A. x ≤ –8 B. x ≥ –5 C. x ≤ –5 D. x ≥ –8

OpenStudy (ad7):

B

Mehek (mehek14):

First you would want to move 6 to the other side so -4 - 6 = ?

OpenStudy (linstead_cpd_freak):

-4 - 6 = -10

alones (alones):

@ad7 Welcome to Openstudy, but here were not allowed to give direct asnwers since it's again ConC.

Mehek (mehek14):

alright so we have 2x ≥ - 10 now we need to isolate x so -10/2 = ?

OpenStudy (linstead_cpd_freak):

-5

Mehek (mehek14):

so x ≥ -5

OpenStudy (linstead_cpd_freak):

Tank you!

Mehek (mehek14):

You're welcome :D

OpenStudy (ad7):

@AloneS Sorry, won't happen again.